OpenStudy (yacoub1993):

Solve the system of equations using matrices. Use Gauss-Jordan elimination. 3x - 7 - 7z = 7 6x + 4y - 3z = 67 -6x - 3y + z = -62

OpenStudy (yacoub1993):

@da_ScienceMan

OpenStudy (yacoub1993):

this is similar to that question @da_ScienceMan

OpenStudy (yacoub1993):

just tell me how to start and i will try to solve it

OpenStudy (da_scienceman):

ok hi. Did u try to go through the other one? and did u verify the steps? I was really in a rush!

OpenStudy (yacoub1993):

i did with my cousin but i didnt get it very clear... but i will try my best to solve this one and see if i can do it myself

OpenStudy (yacoub1993):

just tell me how to start of

OpenStudy (yacoub1993):

@da_ScienceMan u der

OpenStudy (da_scienceman):

yes it is almost like the method yesterday....just begin by writing down the augmented matrix.

OpenStudy (yacoub1993):

i mean wxyz that way

OpenStudy (da_scienceman):

\[\left[\begin{matrix}3 & -7 & -7 | 7 \\ 6 & 4 & -3 | 67 \\ -6 & -3 & 1 | -62 \end{matrix}\right]\]

OpenStudy (da_scienceman):

so have u written d the matrix i just typed? we can proceed...this one is shorter than the one we had yesterday!

OpenStudy (da_scienceman):

@Yacoub1993 u there?

OpenStudy (yacoub1993):

yes i am there, solving the thread

OpenStudy (da_scienceman):

ok so we begin by multiplying row 1 by 2 and adding to row 3. But we keep our row 1 the same as the old matrix. Only row 3 is permitted to change! Next, we multiply row 1 by -2 and add to row 2. Only elements in row 2 are permitted to change. All other rows from the previous step remain unchanged!

OpenStudy (yacoub1993):

i coulnt do it @da_ScienceMan

OpenStudy (yacoub1993):

Solve the system of equations using matrices. Use Gauss-Jordan elimination. 3x - 7 - 7z = 7 6x + 4y - 3z = 67 -6x - 3y + z = -62 A. {( 7, 1, 7)} B. {( 14, 7, -7)} C. {( -7, 7, 14)} D. {( 7, 7, 1)}

OpenStudy (da_scienceman):

ok why do u have the multiple choices? are those answers or what?

OpenStudy (da_scienceman):

@Yacoub1993 brb!

OpenStudy (da_scienceman):

ok so if we begin i can discuss and tell u which values u ll get so u write them down and we solve togewhther. It will make u become an expert in these type of equations. ready?

OpenStudy (da_scienceman):

ok sorry lets knw when u are on @Yacoub93

OpenStudy (yacoub1993):

Anyone help me Please

OpenStudy (yacoub1993):

@da_ScienceMan

OpenStudy (anonymous):

Do you have a TI-83 + calculator? There's a way to solve it on a calculator.

OpenStudy (yacoub1993):

i have this one fx-991ES PLUS

OpenStudy (yacoub1993):

is it similar @AngelicaLS

OpenStudy (anonymous):

I've never heard of that one but do you see a button that says 2nd? If you hit that and the button underneath Matrix which is x^-1, it should take you to a menu.

OpenStudy (yacoub1993):

matA or matB or matC

OpenStudy (yacoub1993):

@AngelicaLS

OpenStudy (anonymous):

doesn't matter. go with mat a. once you get to that menu, do you see dimensions? If so, change them to 3x4 for this problem.

OpenStudy (yacoub1993):

ok

OpenStudy (yacoub1993):

there is no 3x4

OpenStudy (yacoub1993):

maximum is 3x3 @AngelicaLS

OpenStudy (anonymous):

that's strange. let me try something on my calculator. one moment.

OpenStudy (anonymous):

yeah, i don't know why your calculator doesn't allow dimensions larger than 3x3 so i can't teach you how to do it on yours. i do know that the answer is D (7,7,1). hopefully you can use that as a guide to find out the work behind it for yourself. :/